# STaRK: Benchmarking LLM Retrieval on Textual and Relational Knowledge Bases

### v1

@@ 6 Acknowledgement
We thank group members in Jure Leskovec and James Zou’s lab for providing valuable suggestions. We also acknowledge the support of Amazon, DARPA under Nos. N660011924033 (MCS); NSF under Nos. OAC-1835598 (CINES), CCF-1918940 (Expeditions), DMS-2327709 (IHBEM); Stanford Data Applications Initiative, Wu Tsai Neurosciences Institute, Chan Zuckerberg Initiative, Genentech, GSK, Hitachi, Juniper Networks, KDDI, and UCB.
Appendix A Benchmark details
@@
A.1 STaRK-MAG
@@ A.2 STaRK-Prime
Relational tuples.
[(anatomy, expression absent, gene/protein),
(anatomy, expression present, gene/protein),
(anatomy, parent-child, anatomy),
(biological_process, interacts with, exposure),
(biological_process, interacts with, gene/protein),
(biological_process, parent-child, biological_process),
(cellular_component, interacts with, exposure),
(cellular_component, interacts with, gene/protein),
(cellular_component, parent-child, cellular_component),
(disease, associated with, gene/protein),
(disease, contraindication, drug),
(disease, indication, drug),
(disease, linked to, exposure),
(disease, off-label use, drug),
(disease, parent-child, disease),
(disease, phenotype absent, effect/phenotype),
(disease, phenotype present, effect/phenotype),
(drug, carrier, gene/protein),
(drug, contraindication, disease),
(drug, enzyme, gene/protein),
(drug, indication, disease),
(drug, off-label use, disease),
(drug, side effect, effect/phenotype),
(drug, synergistic interaction, drug),
(drug, target, gene/protein),
(drug, transporter, gene/protein),
(effect/phenotype, associated with, gene/protein),
(effect/phenotype, parent-child, effect/phenotype),
(effect/phenotype, phenotype absent, disease),
(effect/phenotype, phenotype present, disease),
(effect/phenotype, side effect, drug),
(exposure, interacts with, biological_process),
(exposure, interacts with, cellular_component),
(exposure, interacts with, gene/protein),
(exposure, interacts with, molecular_function),
(exposure, linked to, disease),
(exposure, parent-child, exposure),
(gene/protein, associated with, disease),
(gene/protein, associated with, effect/phenotype),
(gene/protein, carrier, drug),
(gene/protein, enzyme, drug),
(gene/protein, expression absent, anatomy),
(gene/protein, expression present, anatomy),
(gene/protein, interacts with, biological_process),
(gene/protein, interacts with, cellular_component),
(gene/protein, interacts with, exposure),
(gene/protein, interacts with, molecular_function),
(gene/protein, interacts with, pathway),
(gene/protein, ppi, gene/protein),
(gene/protein, target, drug),
(gene/protein, transporter, drug),
(molecular_function, interacts with, exposure),
(molecular_function, interacts with, gene/protein),
(molecular_function, parent-child, molecular_function),
(pathway, interacts with, gene/protein),
(pathway, parent-child, pathway)]
@@
    "What is the drug that some genes or proteins act as an enzyme upon, where the genes or proteins are expressed in <anatomy>?",
@@
(drug → [indication] → disease → [indication] → drug) & (drug → [synergistic interaction] → drug):
@@
(pathway → [interacts with] → gene/protein → [interacts with] → pathway) & (pathway → [parent-child] → pathway):
@@
(gene/protein → [associated with] → disease → [associated with] → gene/protein) & (gene/protein → [ppi] → gene/protein):
@@
(gene/protein → [associated with] → effect/phenotype → [associated with] → gene/protein) & (gene/protein → [ppi] → gene/protein):
@@
Appendix B Prompts
B.1 Extracting textual requirements
@@
B.2 Combining relational and textual requirements
@@
B.3 Filtering additional answers

### v2

@@ 6 Acknowledgement
We thank group members group members in Leskovec and Zou labs for providing valuable suggestions and conducting benchmark construction. We express our gratitude to the following individuals for their assistance in generating the human-generated queries (ordered by last name):
Michael Bereket, Charlotte Bunne, Yiqun Chen, Ian Covert, Alejandro Dobles, Teddy Ganea, Bryan He, Mika Sarkin Jain, Weixin Liang, Gavin Li, Jiayi Li, Sheng Liu, Michael Moor, Hamed Nilforoshan, Rishi Puri, Rishabh Ranjan, Yanay Rosen, Yangyi Shen, Jake Silberg, Elana Simon, Rok Sosic, Kyle Swanson, Nitya Thakkar, Rahul Thapa, Kevin Wu, Eric Wu, Kailas Vodrahalli.
We especially thank Gavin Li at Stanford University and Zhanghan Wang at New York University for helping build the interactive interface for our SKBs.
We also gratefully acknowledge the support of DARPA under Nos. N660011924033 (MCS); NSF under Nos. OAC-1835598 (CINES), CCF-1918940 (Expeditions), DMS-2327709 (IHBEM); Stanford Data Applications Initiative, Wu Tsai Neurosciences Institute, Stanford Institute for Human-Centered AI, Chan Zuckerberg Initiative, Amazon, Genentech, GSK, Hitachi, SAP, and UCB.
Appendix A Mathematical Definitions of Shannon Entropy and Type-Token Ratio
Shannon Entropy. Shannon Entropy is a measure of the uncertainty in a set of possible outcomes, quantifying the amount of information or disorder within a dataset. It is defined as follows:
$H(X)=-\sum_{i=1}^{n}p(x_{i})\log p(x_{i})$
where $X$ is the set of possible outcomes, $p(x_{i})$ is the probability of occurrence of the outcome $x_{i}$, and $n$ is the total number of unique outcomes. Higher entropy values indicate greater diversity in the distribution of outcomes.[35]
Type-Token Ratio (TTR). The Type-Token Ratio is a measure of lexical diversity, calculated as the ratio of the number of unique words (types) to the total number of words (tokens) in a text. It is defined as follows:
$\text{TTR}=\frac{V}{N}$
where $V$ is the number of unique words and $N$ is the total number of words in the text. Higher TTR values indicate a higher proportion of unique words, reflecting greater lexical diversity. [41]
Appendix B Benchmark details
@@
B.1 STaRK-Amazon
Basic query templates:
metapath Query template
(brand → product) "Can you list the products made by <brand>?"
(product → product) "Which products are similar to <product>?"
(color → product) "Can you provide a list of products that are available in <color>?"
(category → product) "What products are available in the <category> category?"
The final relational template can be composed of multiple basic templates. For example, ‘(color → product ← brand)’ represents one such combination.
B.2 STaRK-MAG
@@ B.3 STaRK-Prime
    "What is the drug that some genes or proteins act as an enzyme upon,
     where the genes or proteins are expressed in <anatomy>?",
@@
(drug → [indication] → disease → [indication] → drug)
 & (drug → [synergistic interaction] → drug):
@@
(pathway → [interacts with] → gene/protein → [interacts with] → pathway)
    & (pathway → [parent-child] → pathway):
@@
(gene/protein → [associated with] → disease → [associated with] → gene/protein)
    & (gene/protein → [ppi] → gene/protein):
@@
(gene/protein → [associated with] → effect/phenotype → [associated with] → gene/protein)
    & (gene/protein → [ppi] → gene/protein):
@@
Appendix C Prompts and LLM versions for Query Synthesization
We summarize the LLM versions in Table 11. We chose these models based on their cost, how accurate they are, and whether they were the latest model during different parts of the project. While we used different LLMs, we checked each step separately to make sure the good quality in our benchmark datasets.
Table 11: Summary of LLM Versions for Query Synthesization
Step STaRK-Amazon STaRK-MAG STaRK-Prime
Step 2: Extracting textual requirements gpt-3.5-turbo-16k claude-2.0 claude-2.0
Step 3: Combining relational and textual requirements claude-2.0, gpt-4-0125-preview
Step 4: Filtering additional answers claude-2.1, claude-2.0, claude-instant-1.2
C.1 Extracting textual requirements
@@
C.2 Combining relational and textual requirements
@@
C.3 Filtering additional answers
@@
Appendix D Instructions for Generating Queries
For the process of generating queries by human, the participants were given a list of entity IDs that we randomly sampled from the entire entity set. Then, they were asked to follow the following instructions with the support of our built interactive demo at https://huggingface.co/spaces/snap-stanford/SKB-Explorer. {mdframed}[backgroundcolor=gray!10, linecolor=black, linewidth=2pt, innerleftmargin=10pt, innerrightmargin=10pt, innertopmargin=3pt, innerbottommargin=10pt] Task:
1) Given the provided entity ID, review the associated document and any connected entities and multi-hop paths.
2) Find interesting aspects of the entities by examining both their relational structures and the textual information available.
3) Write your queries from these aspects such that the entity can satisfy all of them.
Note:
1) Please do not leak the name of the entity in the query.
2) You can skip some queries if you think the knowledge involved is hard to understand.
3) Feel free to be creative with content of your queries, you can also include additional context. There is NO restriction on how you express the queries. After that, we filtering the ground truth answers manually by human validation.